How do you calculate NPV of an investment that a company is financing? The machine cost 400,000. will pay 4.15% and the payment will be $16,600 a month and 5,000 for yearly maintenance. The life of the machine is 10 years.
Answer #1
Monthly payment =$16,600
Monlty interest rate =4.15%
Period =10 years =10*12=120 Months
Present value of cash flows =Annuity factor at 4.15% for 120 periods =23.91322
Present value of cash flows of 16600=$16000*23.91322=$382611.50
yearsly maintanace =5000 per year
Years interest rate =4.15%*12 =49.80%
Present value of cash flows =Annuity factor at 49.80% for 10 periods =1.9727
Present value of cash flows of 5000=$5000*1.9727=$9863.50
Total Present value of Cash Inflows =$382611.50+$9863.50=$392,475
Net present value of investment Amount in $
Total present value of Cash inflows 392475
Less : Cost of invesment -400000
Net present value -7525
NPV of Invesent is $(7525)
this investment is generating Negitive NPV , it not financially Viable.